# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

### v3

@@ V-D IRS
With the aid of IRS, Chen et al. [171] addressed the joint design of the HBF matrix and the phase-shifting matrix of the IRS elements. The authors proposed the deep unfolded WMMSE-MO network (DU-WMMSE-MO), for IRS-assisted mmWave MIMO-OFDM systems. The trainable parameters include the weights for the precoder, phase configuration of the IRS elements and the combiner, and biases for the precoder and phase-shift matrix of the IRS. The algorithm showed improved performance in IRS phase-shift design, utilizing two loops to jointly update the step-size of the beamforming and IRS matrices. Compared to the Armijo backtracking line search method used in classical WMMSE-MO [172], DU-WMMSE-MO [171] provides more flexibility but with a higher runtime.
An IRS-assisted multiuser MIMO full-duplex system is explored in [173]. The active beamforming matrices at the BS and uplink users, along with the passive beamforming matrix at the IRS, are jointly optimized to maximize the weighted sum rate of the system. A deep unfolding approach is proposed to approximate the stochastic SSCA-based beamforming algorithm, reducing the complexity of matrix inversion while maintaining near-optimal performance.
@@
In [174], Min et al. introduced a two-timescale transmission approach to reduce the signal processing complexity and channel training overhead for the IRS-aided multiuser MISO system. To minimize the average transmit power at the analog precoder while ensuring the individual quality of service (QoS) constraints for the users, a primal-dual decomposition-based two-timescale HBF method is proposed and unfolded. This approach allows for efficient beamforming optimization by separating the optimization into two timescales, improving both computational efficiency and performance.
@@
In [175], a low-complexity IRS-assisted downlink MISO system is designed using an unfolded block CD algorithm to address the joint optimization problem of transmit beamforming and IRS phase shifts. The multi-ratio fractional programming structure of the objective function is handled by decoupling the problem into two subproblems: one for optimizing the beamforming at the BS and another for optimizing the phase shifts of the IRS. This is achieved through the use of Lagrange dual transformation, effectively simplifying the optimization process. The numerical analysis validated that the achieved weighted sum rate is comparable to that of [176] but with reduced computational complexity. Furthermore, the authors in [177], have developed an unfolded network to address the joint beamforming task for an IRS-assisted MISO system. The WMMSE and power iteration algorithm are used as an alternating optimization (AO) strategy to derive the local optimal solution for the active and passive beamforming. The proposed network reduce the computational runtime to approximately 3% of the block CD algorithm in [176].
@@
In [178], an IRS-assisted MISO covert symbiotic radio communication system is explored, and a deep unfolding algorithm based on gradient descent is proposed for beamforming design. The goal is to jointly optimize active and passive beamforming under covert communication constraints. The step-size for updating the active and passive beamformers, and Lagrange multipliers form the trainable set. In contrast, Gangyong et al. in [179] focused purely on passive beamforming and unfolded the Riemannian gradient descent algorithm. They have designed the IRS phase shift matrix by maximizing the sum path gain of the channel. In this network, the only trainable parameter is the step-size associated with each Riemannian gradient descent iteration. The network in [179] is simpler compared to that of [178], due to less number of parameters to train but is specific to optimizing the phase shift of the IRS.
@@ V-F Discussions and Takeaways
Precoding design in massive MIMO and HBF systems involves solving non-convex optimization problems under hardware constraints and channel uncertainty. Popular unfolded networks for precoding design are listed in Table V. For instance, IAIDNN [146] unfolds the WMMSE algorithm and replaces complex operations like matrix inversion with a learnable block. This results in a significant reduction in computational complexity while achieving performance close to traditional WMMSE solutions. Deep unfolding enables the incorporation of domain-specific constraints such as the constant modulus constraint in analog beamformers through tailored activation functions. This approach significantly reduces computational complexity by avoiding matrix inversion and accelerates convergence, making it well-suited for real-time hybrid precoding under practical hardware limitations.
Despite their benefits, current unfolded precoding models often assume fixed system dimensions and ideal CSI, which limits generalization and applications in practical systems. Future work should focus on scalable, adaptive unfolding frameworks that can operate reliably in time-varying, multi-user, and wideband settings. Integration with channel estimation modules and robust designs under CSI uncertainty will be key for practical deployment in next-generation wireless networks. Moreover, applying such unfolded frameworks for precoders in real systems and analyzing the performance in real experiments will be very beneficial.
@@ VI Sensing and Communication
TABLE VI: Summary of typical deep unfolding models for ISAC and their advantages and disadvantages.
Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
Unfolded PGA network [191] Unfolds PGA to maximize the system sum rate while minimizing the sensing beampattern error. Achieves $33.5\%$ higher sum rate, $2.5$ dB lower radar beam pattern error compared with SCA, and $65\%$ reduction in runtime and complexity compared with the conventional PGA. Advantages: Ensure convergence in the challenging ISAC HBF designs, perform well in both single use and MU systems. Disadvantages: The convergence speed is relatively slow, high complexity to compute the gradients.
ISAC-Net [194] Unfolds the denoising AMP method for communication and DFT-based sensing in narrowband OFDM ISAC. Achieve communication performance close to OAMP-Net2 [11] and better sensing performance than classical 2D-DFT algorithm. Advantages: Leverages shared resources to overcome challenges in signal processing without prior knowledge of transmitted signals, enhanced accuracy and security. Disadvantages: Beam squint effect is not considered leading to discrepancies extracting sensing parameters (angle of arrival or angle of departure) across various subcarriers in OFDM.
ADMM-Net for OTFS ISAC [196] Unfolds ADMM to estimate delay-Doppler targets and demodulate OTFS symbols under high mobility. With $2$ targets, ADMM-Net achieve $4$ dB gain in NMSE compared with the classical ADMM and ISTA at SNR=$14$ dB. Advantages: Suitable for vehicular/high-mobility scenarios and robust to delay-Doppler spread. Disadvantages: Real time CSI estimation is not addressed.
Double loop unfolded network [197] Unfolds gradient descent to jointly optimize IRS phase shifts in ISAC system. Minimizes worst case CRB for sensing targets. Advantages: Has low computational complexity and fast convergence (within $15$ layers). Disadvantages: Offline training ($1000$ channel realization, $20$ epochs) may not be feasible for frequently changing environments.
@@ VI-A MIMO ISAC
HBF in MIMO ISAC systems poses additional complexity due to balancing communication rates and sensing accuracy. The authors in [191] tackled this challenge by unfolding the iterative PGA algorithm. Their deep unfolding-based method accelerated convergence and enhanced overall system performance, achieving an optimal balance between communication and sensing objectives. In addition, joint optimization of radar communication beamforming in MIMO system was addressed in [192]. The authors incorporated symbol-level precoding with constructive interference constraints to improve energy efficiency by exploiting multi-user interference. Since, the objective problem is computationally complex due to non-convexity and higher dimensional matrix inversions, a custom iterative algorithm is developed. The iterative algorithm is based on block CD and primal-dual optimization (augmented Lagrangian method). Therefore, the trainable parameter set for training the deep unfolded network include step-sizes to update the primal, dual and auxiliary variable, as well as a correction matrix to approximate matrix inversion. The proposed network reduces computational complexity due to small number of trainable parameters, improve scalability, and support real-time implementation.
The large-scale antenna arrays in mmWave MIMO systems enable sub-meter user localization by providing high angular resolution. However, these methods often incur significant computational complexity, particularly in high dimensional mmWave MIMO settings. To improve this, Fan et al. proposed, a fast and accurate direct localization method in [193]. An ADMM solver is developed to solve the optimization problem with low complexity. The ADMM iterations are unfolded to learn the trainable parameters that include penalty and proximal parameters, and weight coefficient for channel gain between the BS and user. The unfolded network converges faster than classical ADMM and computational time is reduced.
@@ VI-B Modern Multicarrier Modulation for ISAC
A joint passive sensing and communication demodulation for OFDM is addressed in [194]. The ISAC-Net is designed by unfolding two algorithms: a denoising-based AMP algorithm for the communication module and the 2D discrete Fourier transform (DFT) for the sensing module. The step-size to update the channel and DFT matrix form the trainable set to balance the communication demodulation and sensing accuracy. Furthermore, Jingcai et al. proposed an unfolded network (FISTA-Net) for joint target parameter estimation and communication in MIMO-OFDM systems [195]. The fast ISTA is unfolded, with learnable parameters such as step-size for weight matrix, enabling high-precision solutions with low computational complexity. The proposed network converges with $8$ layers as compared to $2000$ iteration in fast ISTA. Also, the network achieves an NMSE of $-20.12$ dB while the classical complex ISTA achieved $-16.8$ dB. The problem of estimating parameters of moving targets in high-speed scenarios using OTFS modulation for ISAC systems is explored in [196]. The ADMM iterations are unfolded (ADMM-Net) to accelerate convergence and enhance sparse recovery of target position. Regarding parameter estimation, ADMM-Net [196] and FISTA-Net [195] reduce NMSE significantly. Also, ISAC-Net [194] and FISTA-Net [195] show faster convergence than iterative baselines.
@@ VI-D Discussions and Takeaways
Deep unfolding offers a structured and interpretable framework for tackling challenging optimization problems in ISAC systems. By unrolling classical algorithms like PGD and WMMSE into NN layers, unfolding enables the joint design of beamformers and waveforms that balance sensing accuracy and communication performance. These approaches reduce computational complexity and provide faster convergence compared to traditional iterative solutions. We summarize typical deep unfolding models for ISAC in Table VI. It is observed that deep unfolding exhibits advantages in constant modulus waveform design, optimization of dual-functional beamformers, and power allocation. Indeed, the deep unfolding technique can effectively handle non-convex constraints or objective functions like SINR, sum rate, and CRB, which are infeasible with black-box models.
@@ VII Decoding in Error Correcting Codes
In [36], the authors proposed the TurboNet architecture, to improve turbo decoding by unfolding the max-log MAP algorithm. Trainable parameters were inserted into the max-log-MAP decoding flow like weights for prior LLRs. Numerical analysis demonstarted that TurboNet outperforms classical turbo decoding based on the max-log-MAP algorithm, particularly at moderate-to-high SNRs. Furthermore, TurboNet has been validated through over the air experiments confirming its robustness to SNR mismatches. Moreover, the authors have extended the work by including joint detection and decoding in [208]. A CGD-based OAMP detector (CG-OAMPNet) for the MIMO-OFDM system is proposed by unfolding EP. The robustness of CG-OAMPNet was further validated through over-the-air experiments, demonstrating practical resilience to environmental variations. In correlated channel scenario, CG-OAMPNet outperforms OAMP-Net2 [11] by performance gain of $\sim 2$ dB at BER$=10^{-3}$. The authors in [209] tackle the decoding challenges of marker codes in insertion and deletion channels without perfect CSI. They proposed a network called FBNet to improve detection performance; the forward-backward algorithm is unfolded for marker codes. The trainable weights are the CSI probabilities for insertion/deletion channels. FBNet performs quite close to the classical forward-backward algorithm but is robust to CSI mismatch. FBNet [209] is based on RNN structure and target synchronization errors (insertions/deletions) without accurate CSI, while TurboNet [36] is a deep feedforward DNN structure and targets substitution errors in turbo codes under AWGN channel.
@@
TABLE VII: Summary of typical deep unfolding models for decoding and their advantages and disadvantages.
Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
LADN [39] Introduces learnable ADMM-based decoder with trainable penalties for binary LDPC decoding. Achieves a block error rate reduction of 10$\sim$ 10$\times$ 10 × at SNR $=2$ dB over the baseline ADMM decoder. Advantages: Has small number of learning parameters and fast convergence. Disadvantages: Limited by code length. Inherent codeword structure not exploited leading to high computational complexity.
TurboNet [36] Unfolds max-log MAP decoding steps for turbo codes with learnable weights. Outperforms the conventional max-log MAP scheme by $0.25-0.5$ dB BER. Advantages: Compatible with 3GPP turbo codes and validated through practical experiment. Disadvantages: Training requires fixed interleaver and careful SNR selection.
Quantized Polar BP [210] Learns BP decoding under 1-bit quantization for polar codes over an AWGN channel. Outperforms standard BP with one-bit quantization by $1.2-2$ dB (block error rate). Advantages: Hardware-friendly and applicable to IoT systems with coarse ADCs. Disadvantages: Less scalable. Require large number of trainable weights assigned to each edge of factor graph for decoding.
ADMM-Net [40] Extends LADN [39] to irregular LDPC codes. Unfolded ADMM, learn different penalty weights for each variable node in decoding. Achieve better BER performance than BP and ADMM penalized decoders for irregular LDPC codes. Advantages: Suitable for irregular code graphs. Disadvantages: The initialization strategy for the trainable parameters assumes prior knowledge of the code structure, which makes it less generalizable.
FBNet [209] Introduce a RNN-based unfolding model based on forward-backward decoding for Marker codes over insertion-deletion channels. Outperforms the classical forward-backward algorithm under imperfect CSI. Advantages: Robust to synchronization errors with flexible sequence handling. Disadvantages: Degraded performance under burst insertion-deletion channels.

### v4

@@ V-D IRS
Chen et al. [31] proposed the deep unfolded WMMSE-MO (DU-WMMSE-MO) network for IRS-assisted mmWave MIMO-OFDM systems, jointly designing the HBF and IRS phase-shifting matrices. Trainable parameters include precoder and IRS phase weights and biases. The method improves IRS phase-shift design by jointly updating step-sizes, offering more flexibility than classical WMMSE-MO with Armijo backtracking line search method [240] but with higher runtime. In [128], the authors studied IRS-assisted multiuser MIMO full-duplex systems, using deep unfolding to approximate the stochastic SSCA beamforming algorithm, reducing matrix inversion complexity while achieving near-optimal weighted sum rate performance.
@@
Min et al. [239] proposed a two-timescale unfolded primal-dual HBF method for IRS-aided multiuser MISO systems to reduce signal processing complexity and channel training overhead. By separating optimization into two timescales, the approach efficiently minimizes transmit power while meeting individual quality of service (QoS) constraints, improving computational efficiency and performance.
@@
In [212], a low-complexity IRS-assisted downlink MISO system is designed using an unfolded block CD algorithm, decoupling the joint optimization of transmit beamforming and IRS phase shifts into two subproblems via Lagrange dual transformation. The numerical analysis validated that the achieved weighted sum rate is comparable to that of classical approach in [62] but with reduced computational complexity. In [97], an unfolded network combining WMMSE and power iteration was proposed for joint beamforming, reducing computational runtime to about 3% of the classical approach in [62].
@@
In [78], a deep unfolding algorithm based on gradient descent is proposed for IRS-assisted MISO covert symbiotic radio systems, jointly optimizing active and passive beamforming under covert constraints, with trainable step-sizes and Lagrange multipliers. In contrast, Gangyong et al. in [247] focused only on passive beamforming, using unfolded Riemannian gradient descent to optimize the IRS phase shift matrix by maximizing the channel path gain. The network in [247] is simpler compared to that of [78], due to less number of parameters to train but is specific to optimizing the phase shift of the IRS.
@@ V-E Terahertz
In THz communication systems, ultra-large antenna arrays will be employed to form narrow beams and mitigate severe path loss. This objective naturally leads to the combination of massive MIMO and wideband THz, which introduces significant scalability challenges. In particular, deep unfolding algorithms for precoding must efficiently scale with the large number of antennas and multiple users, whereas existing designs have primarily focused on point-to-point massive MIMO [152] or single-antenna users [209]. Moreover, the frequency selectivity of THz channels makes the design and optimization of beamformers (both analog and digital) more challenging, as they must operate across a large number of subcarriers in wideband systems. The computational complexity of deep unfolding generally grows linearly with the number of subcarriers [152, 209]. This increase in complexity raises concerns about scalability in practical deployments. Moreover, THz systems face hardware constraints such as phase noise, low-resolution ADCs, and synchronization errors, which remain largely unexplored in current deep unfolding designs and present opportunities for further optimization.
TABLE V: Comparison of Classical vs Unfolded methods for Precoder design.
$T$: Layers, $I$: Outer iterations, $J$: Inner iterations, $I_{\text{net}}$: Iterations of network, $N_{RF}$: No. of RF chains, $N_{t}$: No. of transmit antenna, $N_{r}$: No. of receive antenna, $K$: No. of users, $M$: Symbol size, $N_{f}$: No. of subcarriers.
Unfolded Computational Complexity Number of Parameters (Unfolded) Convergence speed (Number of layers)
Model Classical Unfolded Simulation Setup Classical Unfolded
IAIDNN [83] WMMSE:  $\mathcal{O}(N_{RF}^{3})$ $\mathcal{O}(T(K^{2}N_{t}N_{r}^{2}+K^{2}N_{t}^{2}N_{r}+KN_{t}^{2.37}+KN_{r}^{3})$ $TK(3N_{r}^{2}+3M^{2}+MN_{r})+(T-1)K(3N_{t}^{2}+MN_{t})$ Massive MIMO, $N_{t}=128$, $N_{r}=2$ and $K=30$ $100$ iterations 77 Layers
PGA-Net [107] PGA:  $\mathcal{O}(IN_{f}N_{t}(KN_{RF}+N_{t}(K+N_{RF})))$ $\mathcal{O}(TN_{f}N_{t}(KN_{RF}+N_{t}(K+N_{RF})))$ $T(N_{f}+1)$ Single cell downlink MIMO with Rayleigh fading ($N_{t}=12$, $N_{r}=1$, $K=6$ $N_{f}=8$, $N_{RF}=10$) $100$ iterations 55 Layers
ManNet [152] MO-AltMin algorithm:  $I\mathcal{O}(N_{t}N_{f}N_{RF}^{2}+J(3N_{t}N_{RF}+2N_{f}(N_{RF}^{2}+N_{RF})M))$ $(I_{\text{net}}-1)\mathcal{O}(N_{t}N_{f}N_{RF}^{2})+\mathcal{O}(N_{t}N_{f}N_{RF})+I_{\text{net}}\mathcal{O}(2N_{f}N_{RF}^{2}M+T(3N_{t}N_{RF}+2N_{f}N_{RF}M)$ $4TI_{\text{net}}N_{t}$ THz massive MIMO, $N_{t}=\{16,32,64,128\}$, $N_{r}=N_{RF}=M=2$ and $N_{f}=128$ $500$ iterations $30$ Layers
@@ V-F Discussions and Takeaways
Precoding design in massive MIMO and HBF systems involves solving non-convex optimization problems under hardware constraints and channel uncertainty. Table V lists some popular unfolded networks for precoding design and their performance metrics. Deep unfolding enables the incorporation of domain-specific constraints such as the constant modulus constraint in analog beamformers through tailored activation functions. This approach significantly reduces computational complexity by avoiding matrix inversion and accelerates convergence, making it well-suited for real-time hybrid precoding under practical hardware limitations.
Despite their benefits, current unfolded precoding models often assume fixed system dimensions and ideal CSI, which limits generalization and applications in practical systems. Future work should focus on scalable, adaptive unfolding frameworks that can operate reliably in time-varying, multi-user, and wideband settings. Integration with channel estimation modules and robust designs under CSI uncertainty will be key for practical deployment in next-generation wireless networks. Moreover, applying such unfolded frameworks in real systems and analyzing the performance in real experiments will be very beneficial.
On the other hand, more emphasis should be placed on ensuring the scalability of deep unfolded precoding algorithms, especially for massive MIMO wideband systems. This requires not only unrolling conventional iterative algorithms, but also rethinking the unfolding architecture and optimization so that the models scale effectively with the array size, bandwidth, and number of users. Possible solutions include designing unfolding networks with sparse inter-layer connections [152] and employing adaptive pruning or approximated computations to reduce redundant operations across layers [143, 10]. For distributed systems, scalability can also be achieved through hierarchical unfolding architectures that offload the main computational tasks to a central unit, as further discussed in Section X-B3.
@@ VI-A MIMO ISAC
HBF in MIMO ISAC systems introduces complexity by balancing communication rates and sensing accuracy. The authors in [153] addressed this by unfolding the iterative PGA algorithm, accelerating convergence and improving system performance. Similarly, the authors in [231] tackled joint radar-communication beamforming using symbol-level precoding with constructive interference constraints to enhance energy efficiency. To handle the non-convex and high-dimensional optimization, a custom iterative algorithm based on block CD and primal-dual optimization was proposed. The deep unfolded network trains step-sizes and a correction matrix for matrix inversion, reducing complexity, improving scalability, and enabling real-time implementation.
In mmWave MIMO systems, large antenna arrays allow sub-meter localization but with high computational cost. Fan et al. [50] proposed a fast direct localization method using an unfolded ADMM solver with trainable penalty, proximal parameters, and channel gain weights. Also in [232], the authors have unfolded ADMM to design a transceiver for ISAC in a cluttered environment. The learnable parameters consist of the per layer step sizes in the network and the ADMM penalty coefficients. Both these approaches [50, 232] achieve faster convergence and lower computational time than classical ADMM.
@@ VI-B Modern Multicarrier Modulation for ISAC
A joint passive sensing and communication demodulation for OFDM is addressed in [94]. The ISAC-Net is designed by unfolding two algorithms: a denoising-based AMP algorithm for the communication module and the 2D discrete Fourier transform (DFT) for the sensing module. The step-size to update the channel and DFT matrix form the trainable set to balance the communication demodulation and sensing accuracy. Furthermore, Jingcai et al. proposed an unfolded network (FISTA-Net) for joint target parameter estimation and communication in MIMO-OFDM systems [115]. The fast ISTA is unfolded, with learnable parameters such as step-size for weight matrix, enabling high-precision solutions with low computational complexity. The proposed network converges with $8$ layers as compared to $2000$ iteration in fast ISTA. Also, the network achieves an NMSE of 20.12-20.12 dB while the classical complex ISTA achieved 16.8-16.8 dB. The problem of estimating parameters of moving targets in high-speed scenarios using OTFS modulation for ISAC systems is explored in [117]. The ADMM iterations are unfolded (ADMM-Net) to accelerate convergence and enhance sparse recovery of target position. Regarding parameter estimation, ADMM-Net [117] and FISTA-Net [115] reduce NMSE significantly. Also, ISAC-Net [94] and FISTA-Net [115] show faster convergence than iterative baselines.
@@ VI-C IRS-Assisted ISAC
TABLE VI: Comparison of Classical vs Unfolded methods for ISAC.
II: Inner Layer, $J$: Outer Layer, $N_{RF}$: No. of RF chains, $N_{t}$: No. of transmit antenna, $N_{r}$: No. of receive antenna, $N_{b}$: No. of BS antenna, $G$: No. of uniform grid location, $B$: No. of BS, $L_{b}$: No. of uniform grid of angle, $M$: No. of OFDM symbols, $N$: No. of subcarriers, $P$: No. of multipaths.
Unfolded Computational Complexity Number of Parameters (Unfolded) Convergence speed (Number of layers)
Model Classical Unfolded Simulation Setup Classical Unfolded
Unfolded PGA [153] PGA: $IJ\mathcal{O}(4N_{\text{RF}}N_{t}^{2}+3N_{t})$ $\mathcal{O}(IJN_{t}^{2}N_{r})$ $2IJ$ Massive MIMO, $N_{t}=64$, $N_{r}=$, $N_{\text{RF}}=4$. $\sim 500$ iterations $40$ Layers
Unfolded ADMM for localization [50] Direct Source Localization [58]: $\mathcal{O}((GB+\sum_{b=1}^{B}L_{b})^{3.5}\sum_{b=1}^{B}N_{b})$ $\mathcal{O}(((BG)^{2}+(\sum_{b=1}^{B}L_{b})^{2})\sum_{b=1}^{B}N_{b})$ $3J+B$ 5G 3GPP Urban Macro MIMO mmWave channel, $B=4$, $N_{b}=50$, $G=900$, $L_{b}=100$. $30$ iterations 55 Layers
ISAC-Net [94] 2D FFT and cyclic cross correlation [207]: $\mathcal{O}(J(M^{2}+N^{2}))$ $\mathcal{O}(TP(N_{t}N_{r}+MN+N\log(N)+M\log(M)+2N_{t}N_{r}(MN)^{2}))$ $5J$ MIMO OFDM ($N_{t}=N_{r}=8$), 16-QAM modulation, $M=256$, $N=1024$. $500$ iterations 55 Layers
@@ VI-D Discussions and Takeaways
Deep unfolding offers a structured and interpretable framework for tackling challenging optimization problems in ISAC systems. By unrolling classical algorithms like PGD and WMMSE into NN layers, unfolding enables the joint design of beamformers and waveforms that balance sensing accuracy and communication performance. These approaches reduce computational complexity and provide faster convergence compared to traditional iterative solutions. We summarize typical deep unfolding models for ISAC in Table VI and present a comparison in terms of computational metrics. It is observed that deep unfolding exhibits advantages in constant modulus waveform design, optimization of dual-functional beamformers, and power allocation. Indeed, the deep unfolding technique can effectively handle non-convex constraints or objective functions like SINR, sum rate, and CRB, which are infeasible with black-box models.
@@ VII Decoding in Error Correcting Codes
[…]
@@
TABLE VII: Comparison of Classical vs Unfolded methods for Decoding in Error Correcting Codes.
$T$: Layers, $I$: Iterations, $n$: No. of variable nodes, $\Gamma_{a}$: No. of auxiliary variables, $m$: No. of check nodes, $d_{\text{max}}$: Maximum degree of check nodes, $N_{t}$: No. of transmit antenna, $N_{r}$: No. of receive antenna, $t$: Time slot.
Unfolded Computational Complexity Number of Parameters (Unfolded) Convergence speed (Number of layers)
Model Classical Unfolded Simulation Setup Classical Unfolded
ADMM-DL Network [64] ML decoder: $\mathcal{O}(2^{n})$ $\mathcal{O}(n+\Gamma_{a})$ $T(n+\Gamma_{a}+1)$ Irregular LDPC code $[n=2048,m=512]$ 3GPP protocol and $[n=576,m=288]$ WiMAX protocol, BPSK modulation. $24$ iterations 33 Layers
PDD-Net [129] Standard BP decoder: $\mathcal{O}(md_{\text{max}}^{2})$ $\mathcal{O}(n+m(d_{\text{max}}-3))$ $3T$ LDPC code $[n=256,m=128]$, BPSK modulation. $2\times 10^{4}$ iterations $25$ Layers
JCED Net [180] BP decoder: $\mathcal{O}(I(4md_{\text{max}}+md_{\text{max}}^{2}))$ $\mathcal{O}(T(8N{r}N_{t}^{2}+10N_{r}N_{t}+16N_{t}^{2}t+8N_{r}N_{t}t+(8+2d_{\text{max}})m2^{d_{\text{max}}-1}))$ $4T$ mmWave MIMO ($N_{r}=8$, $N_{t}=4$), regular LDPC code [$n=288$, $m=144$]. $100$ iterations 202520-25 Layers